# Exploring Knowledge Tracing in Tutor-Student Dialogues

### v1

@@ Abstract
Recent advances in large language models (LLMs) have led to the development of artificial intelligence (AI)-powered tutoring chatbots, showing promise in providing broad access to high-quality personalized education. Existing works have primarily studied how to make LLMs follow tutoring principles but not how to model student behavior in dialogues. However, analyzing student dialogue turns can serve as a formative assessment, since open-ended student discourse may indicate their knowledge levels and reveal specific misconceptions. In this work, we present a first attempt at performing knowledge tracing (KT) in tutor-student dialogues. We propose LLM prompting methods to identify the knowledge components/skills involved in each dialogue turn and diagnose whether the student responds correctly to the tutor, and verify the LLM’s effectiveness via an expert human evaluation. We then apply a range of KT methods on the resulting labeled data to track student knowledge levels over an entire dialogue. We conduct experiments on two tutoring dialogue datasets, and show that a novel yet simple LLM-based method, LLMKT, significantly outperforms existing KT methods in predicting student response correctness in dialogues. We perform extensive qualitative analyses to highlight the challenges in dialogue KT and outline multiple avenues for future work.
@@ 1 Introduction
Tutoring, which mostly takes the form of dialogues between tutors and students, has proven to be highly effective at improving student learning outcomes [36]. Recent advances in generative artificial intelligence (AI), especially large language models (LLMs), have led to the development of AI-powered tutoring chatbots, such as Khan Academy’s Khanmigo [22] and Carnegie Learning’s LiveHint [4]. By wrapping LLMs around learning content such as textbooks and practice questions, these automated chatbots can interact with students in real time, potentially reaching a larger number of students and alleviating the burden on human tutors.
@@
To use LLMs as tutors, it is important to i) ensure they behave like human tutors and ii) control the safety of their output. Since most of existing LLM-powered tutoring chatbots are developed by either fine-tuning on tutoring dialogue transcripts or prompt engineering, plenty of works over the last few years have studied how to make LLMs behave according to effective tutoring principles. The work in [53] studies the latent decision-making process behind human tutor utterances and shows that adopting decisions by human experts can benefit LLMs in tutoring. The work in [33] identifies tutor moves when they interact with an LLM-powered simulated student agent and shows that move annotation data can make LLMs become better tutors. The work in [17] generates large amounts of synthetic tutoring data via AI roleplay and shows that fine-tuning on this data, in addition to data collected from human tutors and students, makes LLMs more effective in pedagogy. The work in [14] proposes a way to measure tutor uptake, i.e., acknowledging what students have said, which is another beneficial strategy for LLMs to adopt. The work in [42] attempted to summarize tutor utterances into actionable feedback for students using LLMs, but found that the resulting messages were unable to outperform those written by real human teachers.
@@
One important observation is that existing works have almost exclusively studied the tutor turns in dialogues rather than the student turns. A possible reason is that student turns can be noisy in nature, partially due to students not fully mastering the required knowledge, resulting in errors that are harder to analyze than tutor utterance. Take Google’s LearnLM [17] for example: student dialogue turns in their AI roleplay-based synthetic training data are generated by prompting an LLM to make an error that a real student would make. This error information is then fed to the tutor LLM to help the tutor model identify the error and provide corresponding feedback. However, there are two key limitations to this approach: First, assuming LLMs can behave like real students may not be reliable. As recent work [15] shows, LLMs are ineffective at anticipating, parsing, and following flawed reasoning. Second and more importantly, real open-ended student discourse is highly valuable since it may indicate flaws in student knowledge and even reveal specific misconceptions. Therefore, analyzing student dialogue turns can be a form of formative assessment to help tutors and AI-powered chatbots understand student progress and provide more targeted feedback. It can also potentially improve simulated student agents [21] to make them better at reflecting real student behavior, which can, in turn, improve the training of tutor chatbots through reinforcement learning [18, 43].
@@ 1.1 Contributions
In this work, we propose dialogue knowledge tracing, a novel task, which seeks to analyze student discourse using the knowledge tracing (KT) framework [9]. We enable KT on dialogues via a two-step process: i) we use pre-trained LLMs, specifically GPT-4o [37], to annotate existing dialogues with student correctness and knowledge component (KC) tags using Common Core standards, and ii) we train KT models on the annotated data to estimate student knowledge at the dialogue-turn level. We introduce two novel KT methods that leverage the textual content in dialogues: an LLM-based KT method (LLMKT), which fine-tunes the open-source Llama 3 model [34] for the KT objective, and DKT-Sem, which replaces the embeddings of DKT [41] with semantic text embeddings. Using two existing tutor-student math dialogue datasets, we show that i) our new methods significantly outperform existing KT methods, especially when little training data is available, and ii) all KT methods can accurately predict student correctness given enough training data. We also use qualitative analysis and learning curve visualizations to show that LLMKT learns meaningful knowledge state estimates. Finally, we show that GPT-4o’s dialogue annotations are accurate according to human experts.
We acknowledge many limitations of our work upfront since it is, to the best of our knowledge, the first at attempting to perform KT in tutoring dialogues. First, a “time step” in our KT setup corresponds to a single student turn in the dialogue, which may make knowledge states volatile; this setting differs from the scale of a time step in existing KT methods, which usually corresponds to a student responding to an assessment/practice question. We cannot explore setting an entire dialogue as a time step since no existing open-source tutoring dialogue dataset links students across dialogues. Second, our work is restricted to math since analyzing the KCs in math problems is challenging yet well-studied; it remains to be seen whether our approach of Common Core standard tagging generalizes to other subjects. Third, due to the smaller scale of tutoring dialogue datasets compared to those typically used in standard KT works, our methods need to be tested on larger-scale data. One dataset we use in our experiments contains partial dialogues that are cut short for LLM evaluation purposes [35], which further adds to this issue. Therefore, we will publicly release our code [^1] https://github.com/umass-ml4ed/dialogue-kt so that researchers who own a large amount of dialogue data can experiment on their proprietary data.
@@ 2 Related Work
2.1 Knowledge Tracing
KT [9] is a well-studied task in the student modeling literature. It studies the problem of breaking down student learning into a series of time steps practicing certain KCs and using the correctness of each response step to track student knowledge. Classic Bayesian knowledge tracing methods [39] use a latent binary-valued variable to represent whether a student masters a KC or not. Factor analysis-based methods [6, 40] use features in addition to latent ability parameters in a logistic regression setting to predict student response correctness. Since neural networks became popular, plenty of deep learning-based KT methods were developed. These models have limited interpretability since student knowledge is represented as hidden states in complex, deep neural networks. Most of these methods use long short-term memory networks [19] or variants [41, 48], with other variants coupling them with memory augmentation [58], graph neural networks [56], attention networks [16, 38], or pre-trained word embeddings to leverage question text information [28].
Since LLMs started to find success in natural language applications, plenty of researchers have attempted to use them in some ways for KT. Some use the neural network architecture of LLMs and train on student response data [57]. Others use LLMs to represent questions and responses in certain domains such as language learning [11, 50] and computer science education [24, 27]. More recently, [26] attempted prompting-based KT that directly asks LLMs to represent questions and student knowledge states as text. Despite some success in domains such as math, it remains to be seen whether these models for question response data can be useful in text-heavy KT situations such as dialogues.
2.2 Knowledge Component Assignment and Discovery
There exist a lot of prior works on automated KC tagging and discovery, including the notable Q-matrix method [3], which assigns KCs to questions using student response data. More recently, methods such as [47] also learn latent KCs by training deep learning models on KT data and enforcing priors from pedagogical theory. While these methods effectively learn new KCs, they do not produce textual descriptions of KCs, making them difficult to interpret. Several recent works have used LLMs to assign expert-written KCs to math problems via prompting [31, 55], fine-tuning [46], and reinforcement learning [25]. However, none have attempted assigning KCs in dialogues. A prior work has attempted to tag KCs in dialogue turns [10], although using a manual approach that cannot be scaled.
@@ 2.3 Dialogue Analysis
As summarized above, existing works on dialogue analysis have primarily focused on analyzing tutor behavior. The works in [13, 51] automatically classify the “talk moves” that tutors and students make via fine-tuned LLMs, the work in [14] automatically measures tutors’ “uptake” behavior in dialogues, and the work in [1] uses talk moves in dialogues to predict student performance in post-tests. To the best of our knowledge, few prior works have systematically studied the problems of KC tagging, student behavior analysis, and particularly KT, in dialogues.
3 KT in Dialogues
@@
We now formulate the dialogue KT framework. Our goal is to develop methods that can i) estimate student knowledge on KCs involved in each dialogue turn and ii) use these estimates to predict the correctness of student dialogue turns. Intuitively, we can view tutor-student dialogues as sequences of (formative) assessment questions. Each tutor turn acts as a question related to some learning objectives and one or more KCs. These turns can correspond to many different tutor moves: asking a clarification question, identifying a student error, providing actionable feedback, prompting the student to think in a Socratic way, or offering encouragement [14, 52]. Each student turn, on the other hand, is a correct or incorrect response to the task posed by the tutor, which reflects their thinking process and offers insights into their knowledge state. A pair of tutor-student dialogue turns corresponds to a time step in KT. Therefore, we split the dialogue KT framework into three stages: i) identify the KCs involved in each student dialogue turn, ii) classify whether the student response is correct or incorrect, and iii) apply new and existing KT methods on the resulting data.
@@
Formally, we define each dialogue $d$ to be a sequence of alternating tutor turns, $t$, and student turns, $s$, i.e., $d=(s_{0},t_{1},s_{1},\ldots,t_{M},s_{M})$, where $M$ is the total number of turn pairs for that dialogue and $s_{0}$ is present if the student initiates the dialogue. Each relevant turn pair is assigned a correctness label $y_{j}\in\{0,1\}$, where $j\in\{1,\ldots,M\}$ is the turn pair index, $y_{j}=1$ means that $s_{j}$ is a correct response to $t_{j}$, and $y_{j}=0$ means that it is incorrect. In practice, we also use $y_{j}=\text{na}$ in cases where correctness is not well-defined for $t_{j}$, e.g., the turn is off-topic or the tutor is offering emotional support. Each tutor turn is also associated with a set of KCs, $\mathcal{C}_{j}=\{c_{j1},\ldots,c_{jK}\}$, where $K$ is the number of KCs for that turn. These KCs are related to learning objectives outlined by the tutor, which lists knowledge elements that the student must posses to respond correctly. In this work, we use the Common Core math standards [8] as the KCs since they are i) standard and widely adopted and ii) textual in nature and are interpretable to both humans and LLMs. For irrelevant dialogue turns that we do not consider, $\mathcal{C}_{j}=\emptyset$.
@@
In a dialogue turn, we use $z_{jk}\in\{0,1\}$ to denote the student’s unobserved, binary-valued latent mastery status of the KC, $c_{jk}$, that is involved in the current step. $z_{jk}=1$ means that the student masters KC $k$ in dialogue turn $j$. Intuitively, the student needs to master all KCs involved in a turn to respond correctly to the task posed by the tutor. Formally, the probability of a correct student response at turn $j$ can be written as $P(y_{j}=1|z_{j1},\ldots,z_{jK})$. The latent KC mastery depends on the entire dialogue history, i.e., $P(z_{jk}=1)=\phi(c_{jk},t_{\leq j},s_{<j},y_{<j},\mathcal{C}_{<j})$, where $\phi$ returns the probability of the student mastering KC $c_{jk}$, and subscripts $<j$ and $\leq j$ correspond to all past student and (including the current) tutor dialogue turns, respectively. In the next section, we detail how we parameterize and train this function in practice.
@@ 4 Methodology
We now detail our two primary technical contributions. First, we detail how to prompt pre-trained LLMs to automatically annotate dialogue turns to collect the KC and response correctness labels needed for dialogue knowledge tracing. Second, we propose new KT methods for the dialogue KT task that incorporate the rich textual information in tutor-student dialogues. We show a demonstration of the dialogue KT framework and the LLMKT method in Figure 1.
@@ 4.1 Automated Dialogue Annotation
In order to perform knowledge tracing on dialogues, each student conversation turn must first be annotated with two labels: i) the correctness of the student response, and ii) a list of KCs involved in the turn. While it is possible for human experts to provide these annotations, doing so would be incredibly time consuming and difficult to scale. Following recent works that have had success using LLMs to annotate data in educational tasks [45, 53], we use a state-of-the-art LLM, GPT-4o [37], to perform the annotation. We conduct the two annotation tasks separately since we find that doing so leads to improved accuracy compared to combining them. We use a simple, zero-shot chain-of-thought prompting approach and leave exploration of more advanced prompting approaches, such as few-shot prompting or self-reflection, for future work. To facilitate replication, we provide our prompts in an online Supplementary Material [^2] https://osf.io/873ms?view_only=f32472a6560649e692a4e2ba603ef52c. In Section 7, we conduct a human evaluation with math teachers to show that the GPT-provided annotations are mostly accurate and aligned with human judgement.
@@
We note that by annotating a full dialogue with a single prompt, knowledge tracing in real-time becomes impossible since KCs for turn $j$ are assigned with $s_{j}$ and future information in context. This limitation could be alleviated in practice by annotating dialogues one turn at a time without seeing $s_{j}$ when labeling $C_{j}$. However, since doing so would significantly increase the number of API calls and thus the cost of annotation, and we are not currently studying the real-time setting, we leave exploration of individual turn tagging and real-time applications for future work.

### v2

@@ Abstract
Recent advances in large language models (LLMs) have led to the development of artificial intelligence (AI)-powered tutoring chatbots, showing promise in providing broad access to high-quality personalized education. Existing works have studied how to make LLMs follow tutoring principles, but have not studied broader uses of LLMs for supporting tutoring. Up until now, tracing student knowledge and analyzing misconceptions has been difficult and time-consuming to implement for open-ended dialogue tutoring. In this work, we investigate whether LLMs can be supportive of this task: we first use LLM prompting methods to identify the knowledge components/skills involved in each dialogue turn, i.e., a tutor utterance posing a task or a student utterance that responds to it. We also evaluate whether the student responds correctly to the tutor and verify the LLM’s accuracy using human expert annotations. We then apply a range of knowledge tracing (KT) methods on the resulting labeled data to track student knowledge levels over an entire dialogue. We conduct experiments on two tutoring dialogue datasets, and show that a novel yet simple LLM-based method, LLMKT, significantly outperforms existing KT methods in predicting student response correctness in dialogues. We perform extensive qualitative analyses to highlight the challenges in dialogueKT and outline multiple avenues for future work.
@@ 1 Introduction
Tutoring, often in the form of natural-language dialogues between tutors and students, has proven to effectively improve student learning outcomes, in both human tutoring [40] and dialogue-based intelligent tutoring systems [41]. Recent advances in generative artificial intelligence (AI), especially large language models (LLMs), have led to the development of generative AI-powered tutoring chatbots, such as Khan Academy’s Khanmigo [26] and Carnegie Learning’s LiveHint [5]. By wrapping LLMs around learning content such as textbooks and practice questions, these automated chatbots can interact with students in real time, potentially reaching a larger number of students than human tutors, while greatly reducing authoring efforts relative to previous-generation dialogue tutors [23].
@@
To use LLMs as tutors, it is important to i) ensure they use effective pedagogical strategies and ii) control the safety of their output. Most existing LLM-powered tutoring chatbots are developed by either fine-tuning on actual tutor utterances in tutoring dialogue transcripts or prompt engineering. Along these lines, plenty of works over the last two years have studied how to make LLMs follow effective tutoring principles. The work in [59] studies the latent decision-making process behind human tutor utterances and shows that adopting decisions by human experts can benefit LLMs in tutoring. The work in [36] identifies tutor moves when they interact with an LLM-powered simulated student agent and shows that move annotation data can make LLMs become better tutors. The work in [19] generates large amounts of synthetic tutoring data via AI roleplay and shows that fine-tuning on this data, in addition to data collected from human tutors and students, makes LLMs more effective in pedagogy. The work in [15] proposes a way to measure tutor uptake, i.e., acknowledging what students have said, which is another beneficial strategy for LLMs to adopt. The work in [47] attempted to summarize tutor utterances into actionable feedback for students using LLMs, but found that the resulting messages were unable to outperform those written by real human teachers.
@@
One important observation is that existing works involving LLMs have almost exclusively studied the tutor turns in dialogues rather than the student turns. A possible reason is that student turns can be noisy in nature, partially due to students not fully mastering the required knowledge, resulting in utterances that contain errors, which are harder to analyze than tutor utterances. Take Google’s LearnLM [19] for example: student dialogue turns in their AI roleplay-based synthetic training data are generated by prompting an LLM to make an error that a real student would make. This error information is then fed to the tutor LLM to help the tutor model identify the error and provide corresponding feedback. However, there is a key limitation to this approach: assuming LLMs can behave like real students may not be reliable. As recent work [16] shows, LLMs are ineffective at anticipating, parsing, and following flawed reasoning. More importantly, real open-ended student discourse is highly valuable since it may indicate flaws in student knowledge and even reveal specific misconceptions. Therefore, analyzing student dialogue turns can be a form of formative assessment to help tutors and AI-powered chatbots understand student progress and provide more targeted feedback. It can also potentially improve simulated student agents [25] to make them better at reflecting real student behavior, which can, in turn, improve the training of tutor chatbots through reinforcement learning [48, 21].
@@ 1.1 Contributions
In this work, we propose dialogue knowledge tracing (dialogueKT), a novel task, which analyzes student discourse within the knowledge tracing (KT) framework [45]. We enable KT on dialogues via a two-step process: we i) use proprietary LLMs, specifically GPT-4o [42], to annotate dialogues with student response correctness and knowledge component (KC) tags using Common Core standards, and ii) train KT models on the annotated data to estimate student knowledge at the dialogue-turn level. We introduce a novel LLM-based KT method, LLMKT, that leverages the textual content in dialogues, by fine-tuning the open-source Llama 3 LLM [20] on the KT objective. Using two existing tutor-student math dialogue datasets, we show that i) our new methods significantly outperform existing KT methods, especially when little training data is available, and ii) all KT methods can accurately predict student correctness given enough training data. We also use qualitative analysis and learning curve visualizations to show that LLMKT learns meaningful knowledge state estimates. Finally, we show that GPT-4o’s dialogue annotations are accurate, according to human experts. Overall, by assessing student performance in more detail, we can make it possible to apply KT to LLM-human tutoring dialogues, enabling mastery learning and various other pedagogies and forms of detection afforded by KT [45].
We acknowledge many limitations of our work upfront since it is, to the best of our knowledge, the first at attempting to perform KT in open-ended tutoring dialogues. First, a “time step” in our KT setup corresponds to a single student turn in the dialogue, which is highly granular; we cannot explore setting an entire dialogue as a time step since no existing open-source large-scale tutoring dialogue dataset links students across dialogues. Second, our work is restricted to math since analyzing the KCs in math problems is challenging yet well-studied; it remains to be seen whether our approach of Common Core standard tagging generalizes to other subjects. Third, due to the smaller scale of tutoring dialogue datasets compared to those typically used in standard KT works, our methods need to be tested on larger-scale data. One dataset we use in our experiments contains partial dialogues that are cut short for LLM evaluation purposes [39], which further adds to this issue. Therefore, we publicly release our code [^1] https://github.com/umass-ml4ed/dialogue-kt so that researchers who own a large amount of dialogue data can experiment on their proprietary data.
@@ 2 Related Work
2.1 Knowledge Tracing and Knowledge Component Labeling
Knowledge tracing (KT) [45] is a well-studied task in the student modeling literature. It breaks down student learning into a series of time steps practicing certain KCs and uses the correctness of each response step to track student knowledge. Classic Bayesian KT methods use a latent binary-valued variable to represent whether a student masters a KC or not. Factor analysis-based methods [8, 44] use features in addition to latent ability parameters in a logistic regression setting to predict student response correctness. Since neural networks became popular, plenty of deep learning-based KT methods were developed. These models have limited interpretability since student knowledge is represented as hidden states in complex, deep neural networks. Most of these methods use long short-term memory networks [46] or variants [53], coupled with with memory augmentation [64], graph neural networks [62], attention networks [17, 43], or pre-trained word embeddings in question text [31]. Many recent works have attempted to use LLMs in some ways for KT. Some use the neural network architecture of LLMs and train on student response data [63]. Others use LLMs to represent questions and responses in certain domains such as math [24], language learning [55, 13], and computer science education [27, 30]. More recently, [29] attempted prompting-based KT that directly asks LLMs to represent questions and student knowledge states as text. Despite some success in question response modeling, it remains to be seen whether these models can be useful in text-heavy KT situations such as dialogues.
There exist many prior works on automated KC tagging and discovery, including the Q-matrix method [3], which assigns KCs to questions using student response data. More recently, methods such as [52] also learn latent KCs by training deep learning models on KT data and enforcing priors from pedagogical theory. While these methods effectively learn new KCs, they do not produce textual descriptions of KCs, making them difficult to interpret. Several recent works have used LLMs to assign expert-written KCs to math problems via prompting [34, 61], fine-tuning [51], and reinforcement learning [28]. However, none have attempted assigning KCs in open-ended dialogues. A prior work has attempted to tag KCs in dialogue turns [12], although using a manual approach that cannot be scaled.
@@ 2.2 Dialogue Analysis
Existing works on dialogue analysis using LLMs have primarily focused on analyzing tutor behavior. The works in [14, 56] automatically classify the “talk moves” that tutors and students make via fine-tuned LLMs, and the work in [1] uses talk moves in dialogues to predict student performance in post-tests. In pre-LLM dialogue tutors, AutoTutor used latent semantic analysis to infer whether a student response was correct, and when incorrect, whether misconceptions were involved [4]. Why2-Atlas used deep syntactic analysis to convert a student response to a proof, analyzed the proof for errors represented as lack of knowledge [58], and later used Bayesian networks [38]. The Watson Tutor used ensembles of different classifiers on a range of textual features for correctness evaluation [7] and Glicko-2 [18] for KT.
3 Knowledge Tracing in Dialogues
@@
We now formulate the dialogueKT framework. Our goal is to develop methods that can i) estimate student knowledge on KCs involved in each dialogue turn and ii) use these estimates to predict the correctness of student dialogue turns. Intuitively, we can view tutor-student dialogues as sequences of (formative) assessments. Visualized in Figure 1, some tutor turns can be viewed as posing a task that is related to some learning objectives and one or more KCs. They can reflect different tutor moves: asking a clarification question, identifying a student error, providing actionable feedback, prompting the student to think in a Socratic way, or offering encouragement [57, 15]. Student turns, on the other hand, contain correct or incorrect responses to tasks posed by the tutor, which reflect their thinking process and offer insights into their knowledge state. A pair of tutor-student dialogue turns corresponds to a time step in KT. Therefore, we split the dialogueKT framework into three stages: i) identify the KCs involved in each student dialogue turn, ii) classify whether the student response is correct or incorrect, and iii) apply new and existing KT methods on the resulting data.
@@
Formally, we define each dialogue $d$ to be a sequence of alternating tutor turns, $t$, and student turns, $s$, i.e., $d=(s_{0},t_{1},s_{1},\ldots,t_{M},s_{M})$, where $M$ is the total number of turn pairs for that dialogue and $s_{0}$ is present if the student initiates the dialogue. Each relevant turn pair is assigned a correctness label $y_{j}\in\{0,1\}$, where $j\in\{1,\ldots,M\}$ is the turn pair index; $y_{j}=1$ means that $s_{j}$ is a correct response to $t_{j}$, and $y_{j}=0$ means that it is incorrect. In practice, we also use $y_{j}=\text{na}$ in cases where correctness is not well-defined for $t_{j}$, e.g., the turn is off-topic or the tutor is offering emotional support. Each tutor turn is also associated with a set of KCs, $\mathcal{C}_{j}=\{c_{j1},\ldots,c_{jK}\}$, where $K$ is the number of KCs for that turn. These KCs are related to learning objectives outlined by the tutor, which lists knowledge elements that the student must posses to respond correctly. In this work, we use the Common Core math standards [10] as the KCs since they are i) standard and widely adopted and ii) textual in nature and are interpretable to both humans and LLMs. For irrelevant dialogue turns that we do not consider, $\mathcal{C}_{j}=\emptyset$.
@@
In a dialogue turn, we use $z_{jk}\in\{0,1\}$ to denote the student’s unobserved, binary-valued latent mastery status of the KC, $c_{jk}$, that is involved in the current step; $z_{jk}=1$ means that the student masters KC $k$ in dialogue turn $j$. In our framework, the student’s mastery of all KCs involved in a turn determines whether they will respond correctly to the task posed by the tutor. Formally, the probability of a correct student response at turn $j$ can be written as $P(y_{j}=1|z_{j1},\ldots,z_{jK})$. The latent KC mastery depends on the entire dialogue history, i.e., $P(z_{jk}=1)=\phi(c_{jk},t_{\leq j},s_{<j},y_{<j},\mathcal{C}_{<j})$, where $\phi$ returns the probability of the student mastering KC $c_{jk}$, and subscripts $<j$ and $\leq j$ correspond to all past student and (including the current) tutor dialogue turns, respectively. In the next section, we detail how we parameterize and train this function in practice.
@@ 4 Methodology
We now detail our two primary technical contributions. First, we detail how to prompt pre-trained LLMs to automatically annotate dialogue turns to collect the KC and response correctness labels needed for dialogue knowledge tracing. Second, we propose new KT methods for the dialogueKT task that incorporate the rich textual information in tutor-student dialogues. We visualize the dialogueKT framework and the LLMKT method in Figure 1.
@@ 4.1 Automated Dialogue Annotation
In order to perform knowledge tracing on dialogues, each student conversation turn must first be annotated with two labels: i) the correctness of the student response, and ii) a list of KCs involved in the turn. While it is possible for human experts to provide these annotations, doing so would be incredibly time consuming and difficult to scale. Following recent works that have had success using LLMs to annotate data in educational tasks [59, 50], we use a state-of-the-art LLM, GPT-4o [42], to perform the annotation via simple, zero-shot chain-of-thought (CoT) prompting [^2] We release our prompts in an online supplementary material: https://osf.io/873ms?view_only=f32472a6560649e692a4e2ba603ef52c. We conduct the two annotation tasks separately since we find that doing so leads to improved accuracy compared to combining them. In Section 7, we conduct a human evaluation with math teachers to show that the GPT-provided annotations are mostly accurate and aligned with human judgment.
@@
We note that by annotating a full dialogue with a single prompt, KT in real-time becomes impossible since information from future turns is available when annotating any given turn. This limitation could be alleviated in practice by annotating dialogues one turn at a time; however, such an approach is costly and we leave that for future work.